OpenStudy (anonymous):

You have 1/2 a pie and want to share it with three friends. What fraction of the whole pie does each person get?

OpenStudy (anonymous):

(1/3)(1/2)

OpenStudy (anonymous):

oh ok

OpenStudy (selena2345):

is it dividing or multiplying fractions ?@demi567

OpenStudy (selena2345):

@demi567

OpenStudy (anonymous):

dividing fractions

OpenStudy (selena2345):

2/3 @demi567

OpenStudy (anonymous):

how is 2/3 you are diving 1/2 / 3

OpenStudy (selena2345):

bye @demi567

OpenStudy (selena2345):

1/6 @demi567

OpenStudy (selena2345):

1/2/3=1/6 @demi567

OpenStudy (selena2345):

bye @demi567

OpenStudy (anonymous):

1/3 times 1/2 = 1/6